Determine the actual discharge, theoretical discharge and weir coefficient over a 90° V-notch weir for a head of 180 mm flowing 400 liters in 20 seconds.
Answer #1

Actual Discharge, Qact = 400/20

Qact =20 L/s or 0.02 m3/s​​​​​

Theoritical Discharge, Qth = (8/15).((2g)^0.5).(tan(θ/2)).(H^2.5)

Qth = (8/15).((2x9.81)^0.5).(tan(90/2)).(0.18^2.5)

Theoritical Discharge, Qth = 0.0325 m3/s

Coefficient of Discharge, Cd = Qact/Qth = 0.02/0.0325

Coefficient of Discharge, Cd = 0.615